OpenStudy (anonymous):

Solve. x2 – 10x = –13

OpenStudy (anonymous):

-8 ( x ) = -13 -13 / -8 = x

OpenStudy (anonymous):

oh wait. Let me check. i think i got it wrong.

OpenStudy (asylum15):

\[x^2 -10x +13 = 0\] This is a quadratic, know how to solve?

OpenStudy (anonymous):

@Asylum15 It is? Oh.... no wonder i got it wrong.

OpenStudy (asylum15):

You will need to solve this for two X values, using the formula.

OpenStudy (anonymous):